# October 20, 2025

### v1

@@
To be submitted to Physical Review D
© 2025 . CC BY 4.0 licence.
@@ 1 Introduction
The decay ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$  [^1] In the following the decay of the $\phi$ into two charged kaons is implied. is studied in detail by the LHCb collaboration regarding its branching fraction [4], angular distribution [10], and lepton universality [18]. Due to neutral meson mixing and the decay into a CPC\!P-eigenstate, the decay offers the possibility to measure time-dependent CPC\!P-violation phenomena, as for example discussed in Refs. [39, 40]. As the final state is not self-tagging, many observables, such as $P_{5}^{\prime}$, are only accessible if flavour-tagging [^2] Flavour-tagging refers to the determination of the ${B}^{0}_{s}$ flavour at production. can be performed. In this work, the projected sensitivity to the observables of the decay rate are explored for both the tagged and untagged time-dependent and time-independent angular analyses of ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays at hadron colliders. We also present new observables which exhibit reduced dependence on the hadronic form-factors, similar to the observables $Q_{8}^{-}$ and $Q_{9}$ discussed in Ref. [39]. As a benchmark for the capabilities of a dedicated heavy-flavour experiment, the expected performance of the LHCb Upgrade I is used frequently throughout this work. Our extrapolations show that a flavour-tagged and time-dependent angular analysis is feasible at hadron colliders, in contrast to previous suggestions in the literature [40]. To explore constraints on short-distance new physics from these new angular coefficients, we employ an effective field theory (EFT) framework, in which the effective Hamiltonian is expressed as a set of local dimension-six operators, $\mathcal{O}_{i}$, each multiplied by a corresponding Wilson coefficient, $\mathcal{C}_{i}$, representing the effective coupling strength. Using this framework we show that the angular observables obtained provide additional constraints of the Wilson Coefficients.
@@ 2 Differential Decay Rate
In the special case for the ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays under study, the decay-time is also needed to describe the decay-rate due to time-dependent CPC\!P-violation effects that arise because the final state is a CPC\!P-eigenstate. Following Refs. [39, 41], the spin-summed decay-rate depending on $q^{2}$, the helicity-angles and the decay-time, $t$, can be expressed using the angular coefficients $J_{i}$, $h_{i}$, and $s_{i}$ as
@@
$\hskip-35.00005pt\mathbin{\vbox{\hbox{\vtop{\halign{#\cr\hfil$\scriptstyle+$\hfil\cr\kern-1.29167pt\cr$\scriptscriptstyle({-})$\cr}}}}}([J_{i}({q^{2}})-\zeta_{i}\bar{J}_{i}({q^{2}})]\cos(x\Gamma t)-s_{i}({q^{2}})\sin(x\Gamma t)\Big\}f_{i}(\cos\vartheta_{K},\cos\vartheta_{\ell},\varphi),$
@@
with $I_{i}$ and $\tilde{I}_{i}$ denoting time-integrated angular coefficients, where once again $\tilde{I}_{i}$ is defined in terms of the CPC\!P-conjugate terms, $\bar{I}$, as $\tilde{I}_{i}=\zeta_{i}\bar{I}$. The LHCb collaboration has measured the untagged sum of the time-integrated decay-rates [10], which yields a mixture of CPC\!P-symmetric and asymmetric terms, denoted $S_{i}\propto I_{i}+\bar{I}_{i}$ and $A_{i}\propto I_{i}-\bar{I}_{i}$, respectively. Following Ref. [39] the $S_{i}$ and $A_{i}$ observables can be expressed using the $J_{i}$, $h_{i}$, and $s_{i}$ coefficients as
@@
and has the additional corollary that all $6c$ coefficients are zero. The same relations hold for the $S_{i}^{(t)}$, $W_{i}^{(t)}$, $\mathcal{H}_{i}$, and $\mathcal{Z}_{i}$ observables. Using these approximations simplifies the parameter space, and reduces the number of highly correlated parameters present in the PDF. Requiring that the total decay-rate is normalised imposes an additional constraint on the angular coefficients
@@
From Eq. 4, another interesting observation can be made regarding the CPC\!P-asymmetry $A_{C\!P}$. Due to the quantum mechanical oscillations of the ${B}^{0}_{s}$ and the decay into a CPC\!P eigenstate, $A_{C\!P}$ can be measured from the angular distributions alone. This is possible as both amplitudes $A_{X}$ and $\tilde{A}_{X}$ enter the decay rates for the ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ mesons explicitly. The integral of the individual decay rates for ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ can be written as
@@
The same is not true for a decay into a CPC\!P specific final state such as ${{B}^{0}}\!\rightarrow{{K}^{*0}}(\rightarrow K^{+}\pi^{-}){\mu^{+}\mu^{-}}$, for which $A_{C\!P}$ needs to be determined in an extended maximum-likelihood fit including the signal yields.
@@ 3 Optimised Observables
$\mathcal{M}_{5}^{-}$ $=\frac{h_{5}}{\sqrt{\langle J_{2c}+\bar{J}_{2c}\rangle[\langle J_{3}+\bar{J}_{3}\rangle-2\langle J_{2s}+\bar{J}_{2s}\rangle]}}$ (24)
@@
and the corresponding $_{A}P_{i}^{t}$ which are obtained by exchanging the plus in the numerator with a minus. While the asymmetries are close to zero in the SM, their value can change if for example the imaginary parts of the Wilson coefficients are altered. The symmetric $P^{t}_{i}$ observables in ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ behave similar to the $P_{i}$ for the ${{B}^{0}}\!\rightarrow{{K}^{*0}}{\mu^{+}\mu^{-}}$ decays, introduced in Ref. [42], and are similarly sensitive to variations in the vector coupling. This similarity, however, only holds when a time-dependent analysis is performed, as the final-state of the ${B}^{0}_{s}$ decay is a CPC\!P-eigenstate. A tagged, time-integrated analysis of ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays would yield, for example
@@
from which is can be seen that, due the large values of $x$ in the ${B}^{0}_{s}$ system, the oscillatory parts integrate to a small value. This results in a strong suppression of the sensitivity to underlying short-distance effects. In contrast, in a time-dependent analysis, yielding either the $S^{t}_{5}$ (or $P^{t\prime}_{5}$) observable, results in a sensitivity to short-distance effects similar to that of $S_{5}$ (or $P^{\prime}_{5}$) for ${{B}^{0}}\!\rightarrow{{K}^{*0}}{\mu^{+}\mu^{-}}$. Note in a time-independent analysis, the $P_{i}$ and $_{A}P_{i}$ are obtained similarly as in in Eq. 29, but exchanging the $J_{i}$ with the $I_{i}$. The sensitivity to these and the other angular coefficients are evaluated using realistic pseudoexperiments, as described in the next section.
@@ 4 Realistic Projections using Pseudoexperiments
As decay-time measurements rely on the determination of the ${B}^{0}_{s}$ decay-vertex, which is also typically used for candidate selection, the measured ${B}^{0}_{s}$ decay-time distribution is expected to be morphed and no longer exponential. This phenomenon is described in this work using an acceptance function which is multiplied to the PDF. We derive a decay-time acceptance which closely resembles the acceptance in Ref. [52] using an analytical function of the form
@@
where the decay-time, $t$, and the estimated decay-time uncertainty, $\sigma_{t}$, are used. While the decay-time resolution varies across different experiments, we will again use the typical LHCb performance as a benchmark. In typical LHCb analyses, decay-time resolutions of $40\,\mathrm{fs}$ are achieved [53]. This is taken into account by making the width of the Gaussian in Eq. 35 linearly dependent on the decay-time
@@
Note, this corresponds to a reparameterisation of $\sigma_{t}$ in Eq. 35, not a per-candidate resolution, therefore it cannot induce a Punzi-bias [54]. Usually, smaller decay-times have better decay-time resolution, as for large decay-times the momentum resolution contributes more to the uncertainty budget. We stress that Eq. 36 is based on publicly available data from the LHCb collaboration and only acts as an approximation of the typical time-resolution expected with the LHCb experiment.
An example of the generated decay-time distribution is shown in Fig. 2 for the ${B}^{0}_{s}$ and the ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ candidates. In addition, our simulated tagging decisions are illustrated, where tagged ${B}^{0}_{s}$ refers to a ${B}^{0}_{s}$ tag from either the SS or OS tagging algorithms, where the other tagger either agrees or yields no decision. Candidates without a tagging decision, or with differing decisions between the two taggers, are labelled as “untagged”. The characteristic oscillations of the ${B}^{0}_{s}$ and the ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ decay-rates can be seen in the figure. Fig. 2 shows $10^{5}$ signal candidates, without any background pollution.
@@
Figure 2: Example of the generated decay-time distribution for the ${B}^{0}_{s}$ (left) and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ (right) candidates with an overlay of the corresponding decay-rate. The simulated tagging decision is indicated in the histograms, showing a tag for the ${B}^{0}_{s}$ (blue) or the ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ (red) or no conclusive decision (black). The category tagged ${B}^{0}_{s}$ is made up from
@@
In the fit, the well-measured parameters $\Gamma$, $x$, and $y$ are fixed to the central values in Ref. [58]. An example for a fit to the pseudodata generated with $1.5\times 10^{3}$ signal candidates is displayed in Fig. 3, where the three signal categories and the combinatorial background are shown in blue, red, purple, and orange, respectively. Examples of the fit projections for all $q^{2}$-regions and fit-setups can be found in App. C. The fit is validated using large ensembles of pseudoexperiments with large numbers of ${B}^{0}_{s}$ decay candidates. For the $\mathcal{L}_{\rm int}=23\text{\,fb}^{-1}$ scenario, we observe a bias smaller than $20\,\%$ of the statistical uncertainty in the $\mathcal{H}_{1s}$ and $\mathcal{H}_{1c}$ parameters, which disappears for larger sample sizes. Similarly, in the $\mathcal{L}_{\rm int}=23\text{\,fb}^{-1}$ scenario, some observables exhibit undercoverage at the $10\,\%$ level, an effect not present for larger sample sizes. This undercoverage does not affect the quoted sensitivities, since these are obtained directly from the distribution of fitted values across the pseudoexperiments, rather than the individual fit uncertainties. The bias and undercoverage are smaller in the untagged fit-setup, and also are not present for the large sample-sizes. Figure 4 summarises the observed bias and coverage for ensembles of 1000 pseudoexperiments per configuration, with marker colour indicating the number of generated signal candidates, marker type the $q^{2}$-region, and marker opacity the tagging power. An unbiased measurement is indicated by a marker compatible with zero and an accurate coverage is represented by a marker compatible with unity. A similar summary for the untagged time-dependent and untagged time-integrated measurements are given in App. D.
@@ 5 Results
The resulting sensitivity to the CPC\!P-averaged and CPC\!P-asymmetric observables $S^{t}_{i}$ and $A^{t}_{i}$ are shown in Fig. 5 for the different luminosity scenarios. The extrapolations are compared to the precision of the SM predictions, evaluated using flavio [47] with the local form-factors following Ref. [59] and the non-local form-factors are parameterised according to Ref. [60]. The best precision is obtained for the coefficients which are accessible through the untagged sum of the ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ decay-rate, shown in the left panels of Fig. 5. Observables which are accessible only using flavour-tagging are shown in the right panels. If a dedicated heavy-flavour experiment with a similar performance to the LHCb upgrade I collects approximately $80\times 10^{3}$ ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays, as could be expected by the conclusion of Run 5, the measurement for $S^{t}_{1c}$ will likely be at the $10^{-2}$ level and therefore more precise than the current SM predictions. The other observables accessible in the untagged decay-rate which are related to $J_{i}+\tilde{J_{i}}$ will also be measured with an uncertainty at the $10^{-2}$ level. The observables which are only accessible using flavour-tagging could be measured up to a sensitivity of $5\times 10^{-2}$, depending on the flavour-tagging performance. An initial measurement of these observables, with a data set corresponding to an integrated luminosity of $23\text{\,fb}^{-1}$, could yield an uncertainty at the $1\text{--}4\times 10^{-1}$ level, with the exact value depending on the flavour-tagging performance of the machine. In this context also $P_{5}^{t\prime}$ could be measured for the first time in rare ${B}^{0}_{s}$-decays. After the conclusion of Run 5, the precision on $P_{5}^{t\prime}$ should be similar to that obtained by the current measurements in ${B}^{0}$ decays [8].
Figure 5: Extrapolated sensitivity to the CPC\!P-averages $S^{t}_{i}$, CPC\!P-asymmetries $A^{t}_{i}$, $\mathcal{H}_{i}$, and $\mathcal{Z}_{i}$ in comparison to the uncertainty of SM predictions evaluated using flavio [47] where the local form-factors are taken from Ref. [59], and the non-local form-factors follow the parameterisation of Ref. [60]. The expected range for the sensitivity after the respective luminosity milestones is shown in the shaded area. The observables accessible in the untagged measurement are displayed on the left and the observables which require flavour-tagging are shown on the right. Depending on the assumed tagging-power, the sensitivity to the observables which require flavour-tagging varies.
@@
In order to assess the relative advantages of performing time-dependent tagged vs time-dependent untagged fits, the sensitivities for both are evaluated on the same generated data. In the case of an untagged fit, only the $\mathcal{H}_{i}$ observables along with the $S^{t}_{i}$ ($A^{t}_{i}$) observables with $\zeta_{i}$ =1 ($\zeta_{i}$ =-1) can be extracted. In Fig. 6 the change in sensitivity for the tagged and untagged time-dependent fits relative to the untagged sensitivity are shown for these observables. The gain in sensitivity when using an untagged set-up ranges from zero up to six percent of the statistical uncertainty. The gain observed for the $\mathcal{H}_{i}$ observables when used an untagged approach is generally larger than for the $S^{t}_{i}$ and $A^{t}_{i}$ observables. For the $\mathcal{L}_{\rm int}=50\text{\,fb}^{-1}$ scenario improvements on the sensitivity reduce to the $2\,\%$ level and for the $\mathcal{L}_{\rm int}=300\text{\,fb}^{-1}$ scenario, no improvement in the sensitivity is observed any more. This modest gain in sensitivity is not overly surprising: the tagged set-up defaults to the untagged form in the case where neither tagger produces a decision, and, in the case where the taggers disagree, the resulting weights for the ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ PDFs are similar due to the comparable mistag probabilities of the two taggers. In the time-independent fit setup, the comparison of sensitivity is not straight forward, as the fit-parameter now corresponds to a combination of the $S^{t}_{i}$ or $A^{t}_{i}$ and $\mathcal{H}_{i}$, however it is checked that there is no marked gain in precision on the other variables when fixing the $\mathcal{H}_{i}$ observables in the fit.

### v2

# Sebastian Schmitt1[Uncaptioned image], Amr Elmarassy1[Uncaptioned image], Michele Atzeni1[Uncaptioned image], Eluned Smith1[Uncaptioned image]

@@
Published in Physical Review D 113, 036001
© 2026 . CC BY 4.0 licence.
@@ 1 Introduction
The decay ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$  [^1] In the following the decay of the $\phi$ into two charged kaons is implied. is studied in detail by the LHCb collaboration regarding its branching fraction [4], angular distribution [10], and lepton universality [18]. Due to neutral meson mixing and the decay into a CPC\!P-conjugate state, the decay offers the possibility to measure time-dependent CPC\!P-violation phenomena, as for example discussed in Refs. [39, 40, 41]. As the final state is not self-tagging, many observables, such as $P_{5}^{\prime}$, are only accessible if flavour-tagging [^2] Flavour-tagging refers to the determination of the ${B}^{0}_{s}$ flavour at production. can be performed. In this work, the projected sensitivity to the observables of the decay rate are explored for both the tagged and untagged time-dependent and time-independent angular analyses of ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays at hadron colliders. We also present new observables which exhibit reduced dependence on the hadronic form-factors, similar to the observables $Q_{8}^{-}$ and $Q_{9}$ discussed in Ref. [40]. As a benchmark for the capabilities of a dedicated heavy-flavour experiment, the expected performance of the LHCb upgrade I is used frequently throughout this work. Our extrapolations show that a flavour-tagged and time-dependent angular analysis is feasible at hadron colliders, in contrast to previous suggestions in the literature [41]. To explore constraints on short-distance new physics from these new angular coefficients, we employ an effective field theory (EFT) framework, in which the effective Hamiltonian is expressed as a set of local dimension-six operators, $\mathcal{O}_{i}$, each multiplied by a corresponding Wilson coefficient, $\mathcal{C}_{i}$, representing the effective coupling strength. Using this framework we show that the angular observables obtained provide additional constraints of the Wilson Coefficients.
@@ 2 Differential Decay Rate
In the special case for the ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays under study, the decay-time is also needed to describe the decay-rate due to time-dependent CPC\!P-violation effects that arise because the final state is a CPC\!P-conjugate state. Following Refs. [40, 42], the spin-summed decay-rate depending on $q^{2}$, the helicity-angles and the decay-time, $t$, can be expressed using the angular coefficients $J_{i}$, $h_{i}$, and $s_{i}$ as
@@
$\hskip-35.00005pt\mathbin{\vbox{\hbox{\vtop{\halign{#\cr\hfil$\scriptstyle+$\hfil\cr\kern-1.29167pt\cr$\scriptscriptstyle({-})$\cr}}}}}\left([J_{i}({q^{2}})-\zeta_{i}\bar{J}_{i}({q^{2}})]\cos(x\Gamma t)-s_{i}({q^{2}})\sin(x\Gamma t)\right)\Big\}f_{i}(\cos\vartheta_{K},\cos\vartheta_{\ell},\varphi),$
@@
with $I_{i}$ and $\tilde{I}_{i}$ denoting time-integrated angular coefficients, where once again $\tilde{I}_{i}$ is defined in terms of the CPC\!P-conjugate terms, $\bar{I}$, as $\tilde{I}_{i}=\zeta_{i}\bar{I}_{i}$. The LHCb collaboration has measured the untagged sum of the time-integrated decay-rates [10], which yields a mixture of CPC\!P-symmetric and asymmetric terms, denoted $S_{i}\propto I_{i}+\bar{I}_{i}$ and $A_{i}\propto I_{i}-\bar{I}_{i}$, respectively. Following Ref. [40] the $S_{i}$ and $A_{i}$ observables can be expressed using the $J_{i}$, $h_{i}$, and $s_{i}$ coefficients as
@@
and has the additional corollary that all $6c$ coefficients are zero. The same relations hold for the $S_{i}^{(t)}$, $W_{i}^{(t)}$, $\mathcal{H}_{i}$, and $\mathcal{Z}_{i}$ observables. Using these approximations simplifies the parameter space, and reduces the number of highly correlated parameters present in the analysis. Requiring that the total decay-rate is normalised imposes an additional constraint on the angular coefficients
@@
The CPC\!P-asymmetry $A_{C\!P}$ is defined as the difference between the differential decay rates of the ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ mesons normalised to the CPC\!P-averaged rate. Because the ${B}^{0}_{s}$ meson undergoes quantum mechanical oscillations and the studied decay mode has a CPC\!P-conjugate final state, this asymmetry can be determined solely from the angular distributions. This arises because both the $A_{X}$ and $\tilde{A}_{X}$ decay amplitudes enter explicitly in the decay rates of ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ mesons, as can be seen from Eq. 4. The corresponding expression for the integrated differential decay-rate of the ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ are given by
@@
The same is not true for a decay into a CPC\!P specific final state such as ${{B}^{0}}\!\rightarrow{{K}^{*0}}(\rightarrow K^{+}\pi^{-}){\mu^{+}\mu^{-}}$, for which $A_{C\!P}$ needs to be determined in an extended maximum-likelihood fit including the signal yields. Note that we assume that ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ mesons are produced in equal quantities, which is the case at the LHC [43]. For small production asymmetries, experimental measures can be implemented to mitigate the effect on the observables defined above.
@@ 3 Optimised Observables
$\mathcal{M}_{5}^{-}$ $=\frac{h_{5}}{\sqrt{\langle J_{2c}+\bar{J}_{2c}\rangle[\langle J_{3}+\bar{J}_{3}\rangle-2\langle J_{2s}+\bar{J}_{2s}\rangle]}},$ (24)
@@
and the corresponding $_{A}P_{i}^{t}$ which are obtained by exchanging the plus in the numerator with a minus. While the asymmetries are close to zero in the SM, their value can change if for example the imaginary parts of the Wilson coefficients are altered. The symmetric $P^{t}_{i}$ observables in ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ behave similar to the $P_{i}$ for the ${{B}^{0}}\!\rightarrow{{K}^{*0}}{\mu^{+}\mu^{-}}$ decays, introduced in Ref. [44], and are similarly sensitive to variations in the vector coupling. This similarity, however, only holds when a time-dependent analysis is performed, as the final-state of the ${B}^{0}_{s}$ decay is a CPC\!P-conjugate state. A tagged, time-integrated analysis of ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays would yield, for example
@@
from which is can be seen that, due the large values of $x$ in the ${B}^{0}_{s}$ system, the oscillatory parts integrate to a small value. This results in a strong suppression of the sensitivity to underlying short-distance effects. In contrast, in a time-dependent analysis, yielding either the $S^{t}_{5}$ (or $P^{t\prime}_{5}$) observable, results in a sensitivity to short-distance effects similar to that of $S_{5}$ (or $P^{\prime}_{5}$) for ${{B}^{0}}\!\rightarrow{{K}^{*0}}{\mu^{+}\mu^{-}}$. Note in a time-independent analysis, the $P_{i}$ and $_{A}P_{i}$ are obtained similarly as in Eq. 29, but exchanging the $J_{i}$ with the $I_{i}$. The sensitivity to these and the other angular coefficients are evaluated using realistic pseudoexperiments, as described in the next section.
@@ 4 Realistic Projections using Pseudoexperiments
As decay-time measurements rely on the determination of the ${B}^{0}_{s}$ decay-vertex, which is also typically used for candidate selection, the measured ${B}^{0}_{s}$ decay-time distribution is expected to be morphed and no longer exponential. This phenomenon is described in this work using an acceptance function which is multiplied by the normalised decay-rate. We derive a decay-time acceptance which closely resembles the acceptance in Ref. [54] using an analytical function of the form
@@
where the decay-time, $t$, and the estimated decay-time uncertainty, $\sigma_{t}$, are used. While the decay-time resolution varies across different experiments, we will again use the typical LHCb performance as a benchmark. In typical LHCb analyses, decay-time resolutions of $40\,\mathrm{fs}$ are achieved [55]. Usually, smaller decay-times have better decay-time resolution, as for large decay-times the momentum resolution contributes more to the uncertainty budget. This is taken into account by making the width of the Gaussian in Eq. 35 linearly dependent on the decay-time
@@
Note, this corresponds to a reparameterisation of $\sigma_{t}$ in Eq. 35, not a per-candidate resolution, therefore it cannot induce a Punzi-bias [56]. We stress that Eq. 36 is based on publicly available data from the LHCb collaboration and only acts as an approximation of the typical time-resolution expected with the LHCb experiment.
An example of the generated decay-time distribution is shown in Fig. 2 for the ${B}^{0}_{s}$ and the ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ candidates. In addition, our simulated tagging decisions are illustrated, where tagged ${B}^{0}_{s}$ refers to a ${B}^{0}_{s}$ tag from either the SS or OS tagging algorithms, where the other tagger either agrees or yields no decision. Candidates without a tagging decision, or with differing decisions between the two taggers, are labelled as “untagged”. The characteristic oscillations of the ${B}^{0}_{s}$ and the ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ decay-rates can be seen in the figure, where $10^{5}$ signal candidates are displayed without any background pollution.
@@
Figure 2: Example of the generated decay-time distribution for the ${B}^{0}_{s}$ (left) and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ (right) candidates with an overlay of the corresponding decay-rate. The simulated tagging decision is indicated in the histograms, showing a tag for the ${B}^{0}_{s}$ (blue) or the ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ (red) or no conclusive decision (black). The category tagged ${B}^{0}_{s}$ (${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$) is made up from events, where both taggers agree on the classification of the event. Untagged events and events where the tagging decisions disagree make up the ”untagged” category.
@@
In the fit, the well-measured parameters $\Gamma$, $x$, and $y$ are fixed to the central values in Ref. [60]. An example for a fit to the pseudodata generated with $1.5\times 10^{3}$ signal candidates is displayed in Fig. 3, where the three signal categories and the combinatorial background are shown in blue, red, purple, and orange, respectively. Examples of the fit projections for all $q^{2}$-regions and fit-setups can be found in App. C. The fit is validated using large ensembles of pseudoexperiments with large numbers of ${B}^{0}_{s}$ decay candidates. For the $\mathcal{L}_{\rm int}=23\text{\,fb}^{-1}$ scenario, we observe a bias smaller than $20\,\%$ of the statistical uncertainty in the $\mathcal{H}_{1s}$ and $\mathcal{H}_{1c}$ parameters, which disappears for larger sample sizes. Similarly, in the $\mathcal{L}_{\rm int}=23\text{\,fb}^{-1}$ scenario, some observables exhibit undercoverage at the $10\,\%$ level, an effect not present for larger sample sizes. This undercoverage does not affect the quoted sensitivities, since these are obtained directly from the distribution of fitted values across the pseudoexperiments, rather than the individual fit uncertainties. The bias and undercoverage are smaller in the untagged fit-setup, and also are not present for the large sample-sizes. Figure 4 summarises the observed bias and coverage for ensembles of 1000 pseudoexperiments per configuration, with marker colour indicating the number of generated signal candidates, marker type the $q^{2}$-region, and marker opacity the tagging power. An unbiased measurement is indicated by a marker compatible with zero and an accurate coverage is represented by a marker compatible with unity. A similar summary for the untagged time-dependent and untagged time-integrated measurements are given in App. D. The data analysed in this work are generated using the publicly available code in Ref. [61].
@@ 5 Results
The resulting sensitivity to the CPC\!P-averaged and CPC\!P-asymmetric observables $S^{t}_{i}$ and $A^{t}_{i}$ are shown in Fig. 5 for the different luminosity scenarios. The extrapolations are compared to the precision of the SM predictions, evaluated using flavio [49] with the local form-factors following Ref. [62] and the nonlocal form-factors are parameterised according to Ref. [63]. The best precision is obtained for the coefficients which are accessible through the untagged sum of the ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ decay-rate, shown in the left panels of Fig. 5. Observables which are accessible only using flavour-tagging are shown in the right panels. If a dedicated heavy-flavour experiment with a similar performance to the LHCb upgrade I collects approximately $80\times 10^{3}$ ${{B}^{0}_{s}}\!\rightarrow\phi{\mu^{+}\mu^{-}}$ decays, as could be expected by the conclusion of Run 5, the measurement for $S^{t}_{1c}$ will likely be at the $10^{-2}$ level and therefore more precise than the current SM predictions. The other observables accessible in the untagged decay-rate which are related to $J_{i}+\tilde{J_{i}}$ will also be measured with an uncertainty at the $10^{-2}$ level. The observables which are only accessible using flavour-tagging could be measured up to a sensitivity of $5\times 10^{-2}$, depending on the flavour-tagging performance. An initial measurement of these observables, with a data set corresponding to an integrated luminosity of $23\text{\,fb}^{-1}$, could yield an uncertainty at the $1\text{--}4\times 10^{-1}$ level, with the exact value depending on the flavour-tagging performance of the machine. In this context also $P_{5}^{t\prime}$ could be measured for the first time in rare ${B}^{0}_{s}$-decays. After the conclusion of Run 5, the precision on $P_{5}^{t\prime}$ should be similar to that obtained by the current measurements in ${B}^{0}$ decays [8].
Figure 5: Extrapolated sensitivity to the CPC\!P-averages $S^{t}_{i}$, CPC\!P-asymmetries $A^{t}_{i}$, $\mathcal{H}_{i}$, and $\mathcal{Z}_{i}$ in comparison to the uncertainty of SM predictions evaluated using flavio [49] where the local form-factors are taken from Ref. [62], and the nonlocal form-factors follow the parameterisation of Ref. [63]. The expected range for the sensitivity after the respective luminosity milestones is shown in the shaded area. The observables accessible in the untagged measurement are displayed on the left and the observables which require flavour-tagging are shown on the right. Depending on the assumed tagging-power, the sensitivity to the observables which require flavour-tagging varies.
@@
In order to assess the relative advantages of performing time-dependent tagged vs time-dependent untagged fits, the sensitivities for both are evaluated on the same generated data. In the case of an untagged fit, only the $\mathcal{H}_{i}$ observables along with the $S^{t}_{i}$ ($A^{t}_{i}$) observables with $\zeta_{i}$ =1 ($\zeta_{i}$ =-1) can be extracted. In Fig. 6 the change in sensitivity for the tagged and untagged time-dependent fits relative to the untagged sensitivity are shown for these observables. The gain in sensitivity when using an untagged setup ranges from zero up to six percent of the statistical uncertainty. The gain observed for the $\mathcal{H}_{i}$ observables when using an untagged approach is generally larger than for the $S^{t}_{i}$ and $A^{t}_{i}$ observables. For the $\mathcal{L}_{\rm int}=50\text{\,fb}^{-1}$ scenario improvements on the sensitivity reduce to the $2\,\%$ level and for the $\mathcal{L}_{\rm int}=300\text{\,fb}^{-1}$ scenario, no improvement in the sensitivity is observed any more. This modest gain in sensitivity is not overly surprising: the tagged setup defaults to the untagged form in the case where neither tagger produces a decision, and, in the case where the taggers disagree, the resulting weights for the ${B}^{0}_{s}$ and ${\kern 1.79993pt\overline{\kern-1.79993ptB}}{}^{0}_{s}$ decay-rates are similar due to the comparable mistag probabilities of the two taggers. In the time-independent fit setup, the comparison of sensitivity is not straight forward, as the fit-parameter now corresponds to a combination of the $S^{t}_{i}$ or $A^{t}_{i}$ and $\mathcal{H}_{i}$, however it is checked that there is no marked gain in precision on the other variables when fixing the $\mathcal{H}_{i}$ observables in the fit.